OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

-2+3=-2x+2-4

OpenStudy (callisto):

Not really. −2(x + 3) = −2(x + 1) − 4 −2(x) + (−2)(3) = −2(x) + (−2)(1) - 4 So, the next step is....?

OpenStudy (anonymous):

so what does x=?

OpenStudy (callisto):

Can you solve the equation?

OpenStudy (callisto):

Of course, simplify that mess first..